OpenStudy (anonymous):

if the triangles are similar, which ratio shows the relationship of the area of the larger triangle to the area of the smaller triangle?

OpenStudy (anonymous):

if the triangles are similar, which ratio shows the relationship of the area of the larger triangle to the area of the smaller triangle?

OpenStudy (anonymous):

1:4 36:1 2:1 6:1

OpenStudy (anonymous):

first of all, you have to know the ratio of the two sides of triangles first. Do you know what it is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, so they give you two sides, they want a ratio of the bigger triangle to the smaller one. So you take the side of the bigger triangle and divide by the smaller one. that's your ratio of triangles' sides

OpenStudy (anonymous):

6:1??

OpenStudy (maheshmeghwal9):

formula is:----\[Ar.(\Delta 1)/Ar.(\Delta2)=(side1)^2/(Side2)^2\]

OpenStudy (anonymous):

idk wat dat is??

OpenStudy (maheshmeghwal9):

\[(12)^2/(2)^2=144/4=36/1\] ; i.e; 36:1.

OpenStudy (maheshmeghwal9):

It is the answer:)--- 36:1.

OpenStudy (anonymous):

:) thnxx

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (maheshmeghwal9):

Now u see these!